OpenStudy (anonymous):

Determine if (1/3, 3) is a solution to 3x+8y=23. Please show all of your work

OpenStudy (anonymous):

8y = 23 - 3x y = 23/8 - 3x/8 3 =? 23/8 - 3*(1/3)/8 = 23/8 - 1/8 = 22/8 22/8 ≠ 3 So no, it's not a solution.

OpenStudy (anonymous):

thanks so much!!